qa3tdyrew:

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

Vocaloid:

First divide both sides by -4 |x+5| = 4 When working with an absolute value problem consider the positive and negative solutions x + 5 = 4 x + 5 = -4 solve for x in both cases. You should get 2 solutions.